A body at 300 degree celcius radiates 10 joules/sq.cm -sec.If sun radiates 100000joules/sq.cm-sec,then its temp is

Grade:12

2 Answers

AskIITians Expert Hari Shankar IITD
14 years ago

Energy radiated is proportional to T4 (Stefan-Boltzmann law)

Let E = kT4

Then, For the first body, 10 = k(573)4

(Because we 300 C = 300+273 = 573 K)

Therefore,

k = 10/(573)4 = 9.27 x 10-11

For the case of the sun,

E = 100000.

So,

100000 = (9.27 x 10-11)T4

or, T4 = 100000/(9.27 x 10-11) = 1.08 x 1015

So T = (1.08 x 1015 )1/4 = 5732.65 K

 

Ramesh V
14 years ago

By Stefan-Boltzmann law,

energy of thermal radiation emitted per unit time by a black body of surface area A is given by:

U= kAT4

T2 = 573K and U1=105 ,  U2=10

so for given conditions, U1/U2 = (T1/T2)4

                                          T1 = 5730 K

Temp. of sun = 5457 C

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch